Choose a,b,c,d
. The demand for a monopolists product is given by Q 4 monopolists marginal cost is given by MC Q. The profit-m output for this monopolist is A) 100 C)40 B) 44-44 D) 20 There is a payoff matrix of two firms; their different profit collusion or competition (answer 14-15). s are listed when they choose firm B competition 14. 14 5, 28 collusion 27, 5 19,19 firm A competition collusion
Answer #1

Demand function

Q = 400-4P

Inverse demand function

P = (400-Q)/4

P = 100-0.25Q

Total revenue = P *Q = (100-0.25Q)*Q

TR = 100Q - 0.25Q​​(2)

MR =Marginal revenue = d (TR)/dQ

MR = d(100Q-0.25Q2)/dQ

MR = 100-0.5Q

Now to get profit maximization quantity

MC = MR

2Q = 100-0.5Q

2.5Q = 100

Q = 40

Answer: 40